# Tailored Emotional LLM-Supporter: Enhancing Cultural Sensitivity

### v1

@@ 5 Experimental Setup
In this work, we focus our experiment on the open-sourced LLMs. We evaluate models that are in the 7B-8B parameters range due to their popularity and performance, including: Llama-3.1-8B Touvron et al. (2023); Dubey et al. (2024), Qwen-2.5-7B Team (2024), and Aya-Expanse-8B Dang et al. (2024). Here, we prioritize open-source model adaptations for scientific reproducibility. We use default configurations for generation in our experiments, and all the adaptation strategies are implemented as system prompts.
@@ 5.1 Automatic Evaluations
We perform fine-grained automatic evaluations based on both emotional support and cultural awareness. We use GPT-o3-mini OpenAI (2025) as the LLM judge due to its high ability for reasoning and correlation with human judges Tan et al. (2025); Gu et al. (2024).
@@ 5.2 Human Evaluations
Crowd evaluation. To compare the effectiveness of different strategies, we conducted in-culture crowd-sourced evaluations using Prolific.666https://www.prolific.com/ We recruit two people per culture who are fluent in both English and the matching language of the culture.777Most Chinese cultural data comes from individuals in English-speaking countries who identify with Chinese culture, shaping their distress experiences. We selected crowd evaluators to match this context.
To evaluate the best adaptation strategy, we sampled 30 data points per culture for all 3 examined models (a total of 2160 response evaluations), along with the human response and the model-generated responses for all strategies. For each culture, we ask two evaluators to pick the best response in terms of both emotional supportiveness and cultural awareness, and aggregate their ranks of adaptation strategies.
To evaluate the best model, we sampled another 20 data points from one of the top strategies across all models we used, and recruited three annotators per culture for the best-performing model evaluation. Here, we use the same criteria as the previous evaluation. The labels are decided based on a majority vote. If there is a tie, another in-culture annotator is involved to make the judgment.888This only happens 5% for cultural awareness, 13.33% for emotional supportiveness, indicating that cultural awareness is a less subjective task than emotional supportiveness.
The instructions were given in both English and the native language of the culture. The detailed instructions are in the Appendix E.
Expert evaluation. We collaborated with two psychologists experienced in online emotional support to 1. validate the safety of the adaptation strategies, and 2. assess the utility of adapted LLMs for training psychology students in cross-cultural therapy Benuto et al. (2018). Together, we developed evaluation guidelines (Appendix H.1). For safety, 100 responses (50 by human, 50 by LLMs) were evaluated, with 15 double-annotated for agreement due to the time-intensive nature of assessment. For utility, 100 LLM-generated responses were evaluated. Agreement results are in Appendix H.2.
@@ 6 Results & Discussions
Table 3 presents the automatic evaluation results of emotion supportiveness and cultural awareness metrics, whereas Table 14 (Appendix D) displays the results for overall quality. In general, all models can respond with high fluency and appropriateness.
Showing in Table 3, the winning strategy is +cga by combining culture-informed role-playing, guidelines and explicit cultural signal annotations. +cga consistently outperforms other strategies across different cultures and models.
Compared to standard role-playing as a Redditor (no suffix in Table 3), simple culture-informed role-playing (+culture) does not improve emotional supportiveness consistently (7/12 times scores lower). In fact, +culture performs worse than providing explicit, detailed cultural signals or guidelines aimed at cross-cultural consultation for emotional supportiveness. Overall, the +guided strategy is more prominent in providing winning emotional support; however, the difference between +annotation and +guided is small.
By explicitly incorporating cultural considerations when offering emotional support (e.g., through +guided, +annotation or +cga), models generate better responses compared to +culture in both evaluation dimensions. We also have a similar observation over human evaluation results (details in §6.1).
Our results here highlight simple, culture-informed role-playing of a generic person from a culture is likely insufficient to provide context-dependent, culturally sensitive tasks; task-specific guidance or guiding model’s attention with explicit signals within the context could offer better help.
@@ 6.1 Human (Crowd) Evaluation
Adapted LLMs can outperform humans in providing responses. Overall, our in-culture evaluators select +cga and +annotation as the best strategies across models and cultures (Table 15). Interestingly, the human response is often ranked as the worst or second-worst strategy, 10 of 12 times for emotional supportiveness. This result further indicates that LLMs could provide better non-professional emotional support in online communities, as the average online human responses can be non-helpful, and fail to recognize and address emotional needs.
Human and LLM judge show moderate to strong correlations. We compute the Kendall rank correlation coefficient () between humans’ and LLMs’ ranking of adaptation strategies (five strategies). The detailed results are in Table 16 (Appendix E.2). In general, human-model correlations are moderate to strong for emotional awareness, except for German culture and Qwen’s responses for Jewish culture. Correlations for cultural awareness are weaker and even negative for German culture, despite it being a high-resource culture.
Notably, humans and the LLM judge diverge the most on Llama’s responses for cultural awareness in Kendall’s . However, as shown in Table 15 (Appendix E.2), humans and the LLM judge often agree on the best strategy, suggesting that divergences mainly stem from lower-ranked strategies. This highlights the need to improve LLMs as judges of cultural awareness.
No single winning model for all cultures. In-culture evaluators evaluate 20 responses from the strategies of each model. Aya-Expanse-8B leads in cultural awareness across Arabic, Chinese, and German contexts. In emotional awareness, Llama-3.1-8B performs best for Chinese culture, while Aya-Expanse-8B receives the highest ratings from evaluators from Arabic and German cultures. Jewish evaluators prefer Llama-3.1-8B for both cultural awareness and emotional supportiveness.
Our evaluation shows: 1. human-model agrees the majority of the time on the best adaptation strategy; 2. LLM judges show room for improvement in cultural awareness; 3. no single model excels at emotional support across cultures.
6.2 Additional Analysis
We qualitatively evaluate responses generated by Aya-Expanse-8B, which was preferred by most cultural groups. To analyze the types of emotional support provided, we use Llama-3.1-70B to categorize responses based on the framework from Liu et al. (2021). Figure 3 shows the distribution of emotional support message categories of Aya-Expanse-8B’s responses across cultures for the strategy. Additional histograms for other adaptation strategies are included in Appendix I. Overall, model-generated responses use a mix of reflection of feelings, affirmation, suggestions, and information, which is distinctly different from human support messages online that primarily offer suggestions. This pattern is consistent across cultures and adaptation strategies.
@@ 6.3 Evaluation with Clinical Psychologists
Our quantitative results and discussion are in Appendix H.3; key highlights include:
LLMs maintain balanced tone, outperform human responses for cognitive empathy, and emotional empathy. Overall, experts indicate that responses were neither overly positive (only 2%) nor overly negative (0%) compared to human responses (6% at both extremes). LLMs outperform humans in cognitive empathy (30% versus 4%) and emotional empathy (34% versus 2%) with the “Strongly agree” rating.
LLMs do not escalate nor introduce new emotional distress. Psychologists agree that LLMs’ responses do not introduce new distress or escalate existing negative feelings with aggregated positive rating (“Strongly agree” and “Agree”) 80% of the time, compared to 28% for human responses.
Suitable for training new clinical psychologists. A total of 88% of LLM responses were found to be promising for training psychology students in cross-cultural therapy, 46% were rated as strong examples suitable for direct use, while 42% required minor adjustments.
Open-feedback from experts. Several qualitative observations emerged from the evaluation process. First, while LLMs often generated structured responses that explicitly included culturally sensitive elements, resulting in higher scores across all criteria, human responses are more authentic and spontaneous. This raises key questions about whether users might value emotionally genuine responses over polished yet generic ones in a certain context. Second, the consistently empathetic tone of LLM (e.g., “You are not alone”) raises an ethical question: should users be informed they are interacting with AI rather than a human? Third, LLMs show a strong tendency to affirm users’ perspectives, but might occasionally miss the clinical red flag, highlighting the need for caution in relying on LLMs for emotional support and emphasizing the importance of incorporating clinical judgment and ethical safeguards into system design and deployment.
7 Conclusion
The ability of LLMs to provide culturally sensitive peer emotional support has been largely overlooked in prior research. To address this gap, we introduce CultureCare, a fine-grained, multicultural dataset encompassing four cultures. Using CultureCare, we evaluate three state-of-the-art LLMs with several carefully designed adaptation strategies. Our findings reveal that: 1. Incorporating shallow cultural information is insufficient compared to using specific, contextualized, guideline-aligned cultural adaptations; 2. LLMs have shown potential to provide emotional support that, in many cases, may surpass that of anonymous online interactions. In collaboration with professional psychologists, we explore the potential of culturally adapted LLMs for training psychology students in culturally informed therapy. Experts highlight their potential usefulness. Overall, our results demonstrate the promise of culturally adapted LLMs as effective tools for enhancing emotional support and advancing culturally sensitive counselling training.
@@ 8 Limitations
In this paper, we investigate culturally aware emotional support using data from Reddit. We acknowledge that, the platform, its users and annotations may introduce representational biases, and thus do not offer a comprehensive representation of any particular culture. In future work, we aim to collect a more diverse and representative dataset through alternative sources and extensive large scale annotations.
Our work explicitly focuses on single-turn support as we focus on online peer settings (which is mostly single-turn in nature), and as a first step in validating LLMs’ ability in providing culturally sensitive support. Future work can utilize CultureCare and extend to multi-turn in different settings.
We also attempt to move away from the common practice of using language (which is often determined by data availability) or nationality (which is typically unavailable in anonymous online communities) as a boundary for cultures. While our approach has certain limitations, it also offers a novel contribution by focusing on self-identified cultural identity and on the underlying causes of emotional stress due to cultural factors as expressed by the users themselves.
Our experiments reveal that the generated responses are sometimes easily identifiable as machine-generated due to distinctive patterns and stylistic features (we refer to them as the AI voice). Prior work has highlighted potential safety concerns associated with anthropomorphic AI systems Deshpande et al. (2023); Akbulut et al. (2024); Moore et al. (2025). This raises a safety versus utility design dilemma regarding whether LLMs intended for emotional support should strive to resemble natural human communication more closely or maintain a distinctive AI voice, and even the design choice should be the same for different cultural groups. In this study, we focus on single-turn interactions, where we think the harm is easier to avoid. We recognize that these issues may become more pronounced in multi-turn conversations due to increasing reliance on interpreting contextual cues from earlier turns, and when the model is presented with adversarial text. These challenges are well beyond the scope of this work, and we plan to address them in future research.
@@
The data used in this study are publicly available online, and the demographic information is voluntarily provided in the original posts on Reddit. While this work explores promising strategies for automatically providing culturally aware emotional support, we do not recommend using our method directly without human verification and large-scale robustness and safety testing. This research is approved by the ethics committee of the Technical University of Darmstadt (EK 121/2024).
@@ A.6 Demographic Information
We extract demographic information from posts using the prompt in Figure 5 with GPT-4o-Mini and manually validated the results. We follow a set of demographic attributes from the WVS Haerpfer et al. (2022), which are included in Table 9. In some cases, only high-level cultural indicators, such as “Arabic” or “Chinese” are extractable.
@@ A.7 Cultural Themes on Emotional Distress and Support
| Culture | Themes |
| LGBTQ+ Rejection: came out … as gay … I was not met with acceptance | |
| Gender Roles: …expects a woman to depend fully on a man and her family | |
| Strict Parenting: my dad is extremely strict | |
| Chinese | Mental Health Invalidation: Depression doesn’t exist |
| Family Dynamics: It feels like she only lives her life for me | |
| Strict Parenting: I live with my tiger parents | |
| Financial Stress: I’m on social welfare and feel ashamed | |
| Gender Roles: …men have more responsibilities … may not be … much fun | |
@@ A.8 Types of Cultural Signals and Support Strategies per Culture
Figure 7 reveals the differences between the types of cultural signals among the 4 cultures. Each culture prioritizes certain signals over others, reflecting unique cultural characteristics. Arabic posts emphasize values and tend to state culture-specific demographics more explicitly (e.g., I’m an arab guy from Sudan). Chinese and German posts prioritize norms and morals, indicating a strong focus on shared principles. In contrast, Jewish posts place a significant emphasis on cultural concepts (e.g., rabbi, Shomer Shabbat), with norms and morals playing a secondary but still important role. Across all cultures, signals of types artifacts, knowledge, and language have minimal presence.
The figure also illustrates cultural differences in emotional support strategies in human responses. The most prominent strategy in all cultures is providing suggestions. Beyond the top strategy, Arabic and Chinese, both emphasize self-disclosure and affirmation. German strategies are more balanced, and Jewish culture uses the information strategy more than the 3 other cultures. The least common strategy is asking questions.
@@ A.9 Intensity of Emotional Distress and Empathy of Responses
Table 12 shows the average scores for the intensity of distress messages in posts and the empathy of human responses across cultures and overall. Jewish posts exhibit the highest intensity (2.07), followed closely by German posts (1.98), indicating that these cultures potentially express distress more vividly compared to Arabic and Chinese. In terms of empathy, Arabic responses show the highest level (3.27), followed by Jewish (3.18), suggesting a more explicit (verbal) approach to emotional support in these cultures. In contrast, Chinese responses demonstrate the lowest empathy (2.15), while German responses fall in the middle (2.73), reflecting a more moderate engagement. Overall, the data highlights cultural variability in both the expression of distress and empathetic responsiveness. It is important to note that since annotators evaluated posts and responses from within their own cultures, these scores may reflect culturally internalized norms. In future work, we plan to explore how perceptions shift when out-of-culture annotators assess the same content.

### v2

@@ 5 Experimental Setup
In this work, we focus on open-source LLMs, prioritizing models in the 7B–8B parameter range due to their strong performance and practical deployment cost, making them well-suited for agentic systems. Our primary evaluation includes Llama-3.1-8B Touvron et al. (2023); Dubey et al. (2024), Qwen-2.5-7B Team (2024), and Aya-Expanse-8B Dang et al. (2024). To examine the robustness of our findings, we additionally test the larger variants of these models. We emphasize open-source models to ensure scientific reproducibility. We use default configurations for generation, and all the adaptation strategies are implemented as system prompts.
@@ 5.1 Automatic Evaluations
We perform fine-grained automatic evaluations based on both emotional support quality and cultural awareness. We use GPT-o3-mini OpenAI (2025) as the LLM-as-a-Judge due to its high ability for reasoning and correlation with human judges Tan et al. (2025); Gu et al. (2024).
@@
Language quality metrics are also derived from Liu et al. (2021) and LaFromboise et al. (1991a), measures: 1. Fluency - the response should be coherent and easy to understand; 2. Communication - the response is appropriate.
We evaluate all aspects on a 5-point scale, which is a common setup in prior work Rashkin et al. (2019); Liu et al. (2021). Following the evaluation prompt-generation method in G-Eval Liu et al. (2023), we use ChatGPT to create prompts that consist of step-by-step evaluation guidelines. See Appendix C.2 for all prompts.
@@ 5.2 Human Evaluations
Crowd evaluation. To compare the effectiveness of different strategies and LLMs, we conducted in-culture crowd-sourced evaluations using Prolific.777https://www.prolific.com/ We recruit two people per culture who are fluent in both English and the matching language of the culture.888Most Chinese cultural data comes from individuals in English-speaking countries who identify with Chinese culture, shaping their distress experiences. We selected crowd evaluators to match this context.
Best adaptation strategy. To evaluate the best adaptation strategy, we sampled 30 posts and corresponding responses for every culture and LLM examined. This results in 2880 evaluations in total. For each post, we display both the human and the model-generated responses. For each evaluation instance, we ask two evaluators to pick the best response in terms of both emotional supportiveness and cultural awareness, then aggregate their ranks of adaptation strategies.
Best LLM. Next, to evaluate the best model, we sampled 20 posts and corresponding responses from the best-performing strategy (+cga), and recruited 3 annotators per culture to find out the best-performing LLM. This results in 720 evaluations in total. Here, we use the same criteria as the previous evaluation. The labels are decided based on a majority vote. If there is a tie, another annotator from the same culture is involved to make the judgment.999This only happens 5% for cultural awareness, 13.33% for emotional supportiveness, indicating that cultural awareness is a less subjective task than emotional supportiveness. The instructions were given in both English and the native language of the culture. The detailed instructions are in the Appendix E.
Expert evaluation. We collaborated with two psychologists experienced in online emotional support to validate the safety of the adaptation strategies (including cognitive and emotional empathy) and assess their utility for training psychology students in cross-cultural therapy (Benuto et al. 2018, including cultural understanding and usefulness for teaching cultural competence). Culturally sensitive emotional support is a well-established component of psychology education, and our use of LLMs reflects real needs identified by practitioners. Together, we developed evaluation guidelines (Appendix H.1) and evaluated 200 responses, illustrating one way our work could support real-world training and educational applications.
| Model | Arabic | Chinese | German | Jewish | ||||
| Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | |
| Aya-Expanse-8B | +cga | +cga | +a | +a | +a | +a | +a | +a |
| Qwen-2.5-7B | +g | +cga | +a | +cga | +cga | +cga | +a | +cga |
| Llama-3.1-8B | +cga | +cga | +g | +cga | +g | +a | +cga | +cga |
@@ 6 Results & Discussions
Table 4 shows the automatic evaluation results for both the emotion supportiveness and cultural awareness aspects. Across models, the strategy showing the best culturally-aware responses (i.e., blue boxes) is +cga, demonstrating the effectiveness of combining culture-informed role-playing, cross-cultural-competence guidelines and explicit cultural signal annotations. Moreover, these results also show that simple culture-informed role-playing alone (+culture) is not enough for offering the most culturally-aware responses. In fact, +culture performs worse than providing explicit, detailed cultural signals (+annotation) or guidelines aimed at cross-cultural consultation (+guidelines) for both emotional supportiveness and cultural awareness. By explicitly incorporating cultural considerations when offering emotional support (e.g., through +guided, +annotation or more significantly through +cga), models generate better responses compared to +culture in both evaluation dimensions. Additionally, we observe similar trends reflected in the human evaluation results (Table 5). More details on the human evaluation are in the next subsection and in §6.1. We also observe consistent patterns on two additional 70B-parameter models (Llama and Qwen, Table 16; Appendix D).
The automatic evaluation results for the language quality of the responses display nearly perfect scores across strategies and models (Table 15), we therefore focus on emotional supportiveness and cultural awareness evaluation in the remainder of this work .
@@ 6.1 Human (Crowd) Evaluation
Adding cultural annotations is essential for more culturally-aware emotional support. As shown in Table 5, across all LLMs and all cultures, the winning strategy for the cultural aspect is a mix of +a and +cga, noting that for both strategies, the cultural signals are included in the input. This shows that high-quality culture annotations can surely enhance the cultural awareness of an emotional support message.
Human and LLM show moderate to strong correlations. We compute the Kendall rank correlation coefficient () between humans’ and LLMs’ ranking of adaptation strategies. For Arabic and Chinese, human-model correlations, averaged over LLMs, are moderate to strong for emotional awareness, namely =0.8 for Arabic and 0.66 for Chinese. For German and Qwen’s responses to the Jewish culture, it is, however, low. Correlations for cultural awareness are weaker (e.g., 0.6 for Arabic and 0.47 for Chinese, averaged over LLMs). However, as shown in Table 5, humans and the LLM-as-a-Judge often agree on the best strategy per model, suggesting that divergences mainly stem from lower-ranked strategies, i.e., humans and LLMs agree on the best but not the worst. The detailed results are in Table 17 (Appendix E.2).
@@ 6.2 Evaluation with Clinical Psychologists
LLMs do not escalate nor introduce new emotional distress. Psychologists agree that LLMs’ responses do not introduce new distress or escalate existing negative feelings with aggregated positive rating (“Strongly agree” and “Agree”) 80% of the time, compared to 28% for Reddit responses.
Adapted responses support training. Psychologists found 88% of LLM responses promising for cross-cultural therapy training, with 46% rated as strong examples and 42% requiring minor adjustments, showing their potential utility in educational contexts.
Qualitative insights. LLMs generate structured, culturally sensitive, and empathetic responses, often including clear guidance, validation, and reassurance. Human replies, while more spontaneous and authentic, sometimes lack consistency or depth in addressing cultural and emotional needs. Experts noted ethical considerations, such as the need for transparency when users interact with AI, the risk of missing clinical red flags, and questions about whether models should closely mimic humans or retain a distinct AI voice. These findings suggest that, in controlled settings, LLMs can complement human support by providing reliable, culturally informed guidance while highlighting areas where human judgment remains essential. 101010We emphasize that these results are specific to our dataset and evaluation context, which involved brief, reactive online peer responses, and do not imply a global superiority of LLMs over human supporters.
7 Conclusion and Future Work
The ability of LLMs to provide culturally sensitive peer emotional support has been largely overlooked. To address this, we introduce CultureCare, a fine-grained dataset spanning four cultures, and evaluate three state-of-the-art LLMs with multiple adaptation strategies. Our results show that shallow cultural cues are insufficient, while contextualized, guideline-aligned adaptations substantially improve performance. Collaborating with professional psychologists, we demonstrate the potential of culturally adapted LLMs for training psychology students. For this emerging research area, we focus on single-turn interactions to establish a reliable foundation for evaluation, with future work extending to multi-turn dialogues, broader cultural contexts, and real-world applications.
@@ 8 Limitations
Data source. In this paper, we investigate culturally aware emotional support using data from Reddit. We acknowledge that the platform, its users and annotations may introduce representational biases, and thus do not offer a comprehensive representation of any particular culture. In future work, we aim to collect a more diverse and representative dataset through alternative sources and extensive large-scale annotations.
Dataset size. Our dataset contains 462 posts, which is smaller than many generic emotion support datasets. However, it includes over 4,000 fine-grained annotations created with in-culture annotators and quality checkers, making it the largest and most detailed resource of its kind. While its size may limit certain large-scale analyses, we believe the dataset’s cultural depth and annotation quality make it highly valuable for evaluating LLMs. We discuss this in detail in Section A.10.
Culture and language. We also attempt to move away from the common practice of using language (which is often determined by data availability) or nationality (which is typically unavailable in anonymous online communities) as a boundary for cultures. While our approach has certain limitations, it also offers a novel contribution by focusing on self-identified cultural identity and on the underlying causes of emotional stress due to cultural factors as expressed by the users themselves.
Cultural coverage and model bias. We acknowledge that models can exhibit cultural biases, and fully addressing these biases across all cultures remains an open challenge. In this work, we explored adaptation for four cultures (Jewish, Arabic, Chinese, and German), showing that culturally informed prompting and the incorporation of explicit cultural signals can improve culturally sensitive emotional-support responses in online settings, though these findings may generalize to other contexts only with careful attention and validation.
@@
All data used in this study are publicly available posts from Reddit, an anonymous forum, ensuring user identities are not disclosed. Our work only collects posts, and the demographic information is voluntarily provided in the original posts on Reddit. While this work explores promising strategies for automatically providing culturally aware emotional support, we do not recommend using our method directly without human verification and large-scale robustness and safety testing. This research is approved by the ethics committee of the Technical University of Darmstadt (EK 121/2024).
@@ Appendix A Additional Information about CultureCare
Figure 2 illustrates our overall dataset creation pipeline.
@@ A.4 Annotation Guidelines
To protect annotators’ mental health, we implemented the following measures: 1. Full transparency about the content, with clear warnings and labels for discussions involving mental distress, abuse, PTSD, and related topics. 2. Allowing annotators to terminate the task at any time.
@@ A.6 Demographic Information
We extracted demographic information from posts using the prompt in Figure 5 with GPT-4o-mini and manually validated the results. We follow a set of demographic attributes from the WVS Haerpfer et al. (2022), which are included in Table 11. In some cases, only high-level cultural indicators, such as “Arabic” or “Chinese” are extractable.
@@ A.8 Types of Cultural Signals and Support Strategies per Culture
Figure 7 reveals the differences between the types of cultural signals among the 4 cultures. Each culture prioritizes certain signals over others, reflecting unique cultural characteristics. Arabic posts emphasize values and tend to state culture-specific demographics more explicitly (e.g., I’m an arab guy from Sudan). Chinese and German posts prioritize norms and morals, indicating a strong focus on shared principles. In contrast, Jewish posts place a significant emphasis on cultural concepts (e.g., rabbi, Shomer Shabbat), with norms and morals playing a secondary but still important role. Across all cultures, signals of the types artifacts, knowledge, and language are minimally present.
The figure also illustrates cultural differences in emotional support strategies in human responses. The most prominent strategy in all cultures is providing suggestions. Beyond the top strategy, Arabic and Chinese, both emphasize self-disclosure and affirmation. German strategies are more balanced, and Jewish culture uses the information strategy more than the three other cultures. The least common strategy is asking questions.
@@ A.9 Intensity of Emotional Distress and Empathy of Responses
Table 13 shows the average scores for the intensity of distress messages in posts and the empathy of human responses across cultures and overall. Jewish posts exhibit the highest intensity (2.07), followed closely by German posts (1.98), indicating that these cultures potentially express distress more vividly compared to Arabic and Chinese. In terms of empathy, Arabic responses show the highest level (3.27), followed by Jewish (3.18), suggesting a more explicit (verbal) approach to emotional support in these cultures. In contrast, Chinese responses demonstrate the lowest empathy (2.15), while German responses fall in the middle (2.73), reflecting a more moderate engagement. Overall, the data highlight cultural variability in both the expression of distress and empathetic responsiveness. It is important to note that since annotators evaluated posts and responses from within their own cultures, these scores may reflect culturally internalized norms. In future work, we plan to explore how perceptions shift when out-of-culture annotators assess the same content.
@@
A.10 Further Discussions on Dataset Size
While our dataset contains 462 posts, we argue that scale alone does not determine the usefulness of an evaluation benchmark. Prior work shows that LLM benchmarks with only a few hundred examples can still provide highly reliable insights. For example, HumanEval (Chen et al., 2021, 164 problems) and DrawBench (Saharia et al., 2022, 200 instances) are widely used in state-of-the-art evaluations. Although smaller in absolute size than generic emotion support datasets, ours includes over 4,000 fine-grained annotations with in-culture annotators and quality checkers. As shown in Tables 1 and 2, it is the largest of its kind and uniquely rich in cultural and emotional details. We further argue that an overemphasis on dataset size risks obscuring contributions that prioritize nuance, cultural specificity, and human well-being; context-rich resources such as ours provide insights that scale alone cannot, and are essential for comprehensive LLM evaluation.